# HalluGuard: Demystifying Data-Driven and Reasoning-Driven Hallucinations in LLMs

### v1

@@ 3.3 Hallucination Quantification via HalluGuard
Let $\mathcal{K}$ denote the NTK Gram matrix with eigenvalues $\lambda_{1}\geq\cdots\geq\lambda_{r}>0$ and condition number $\kappa(\mathcal{K})=\lambda_{\max}/\lambda_{\min}$. Let $J_{t}$ be the step-$t$ input–output Jacobian of the decoder, and define $\sigma_{\max}:=\sup_{t}\|J_{t}\|_{2}$ as the uniform spectral bound(note that $\sigma_{\max}$ is independent of the spectrum of $\mathcal{K}$).
Under Assumptions A1–A3, a standard NTK approximation argument yields $\inf_{u\in U_{h}}\|u^{*}-u\|\;\leq\;C_{d}\,\det(\mathcal{K})^{-c_{d}}\,\|u^{*}\|,$ so that $\det(\mathcal{K})$ capture the representations in systematic bias.
@@
In summary, $\det(\mathcal{K})$ quantifies representational adequacy, $\log\sigma_{\max}$ captures rollout amplification, and $\log\kappa^{2}$ penalizes spectral instability, together forming a compact and tractable proxy consistent with the Hallucination Risk Bound. The lightweight projection layers are self-supervised spectral calibration modules, optimized offline (via AdamW) to align NTK spectral properties across heterogeneous backbones into a stable, comparable geometric space—without hallucination labels or task-specific supervision, with the backbone fully frozen and zero runtime overhead during inference. Detailed proofs are provided in Appendix B.
@@ 4.1 Evaluation Setup
Benchmarks. We evaluate across 10 widely used benchmarks spanning three distinct categories. For data-grounded QA, we include RAGTruth(Niu et al., 2024), NQ-Open(Kwiatkowski et al., 2019), HotpotQA(Yang et al., 2018) and SQuAD(Rajpurkar et al., 2016), which emphasize factual correctness through external evidence. For reasoning-oriented tasks, we use GSM8K(Cobbe et al., 2021), MATH-500(Hendrycks et al., 2021), and BBH(Suzgun et al., 2022), which require multi-step derivations prone to compounding errors. Finally, for instruction-following settings, we consider TruthfulQA(Lin et al., 2022a), HaluEval(Li et al., 2023) and Natural(Wang et al., 2022), which probe hallucinations under open-ended or adversarial prompts.
Baselines. We compare HalluGuard with 11 competitive detectors spanning diverse strategies. Uncertainty-based methods include Perplexity(Ren et al., 2022), Length-Normalized Predictive Entropy(LN-Entropy)(Malinin and Gales, 2020), Semantic Entropy(Kuhn et al., 2023), Energy Score(Liu et al., 2020) and P(true)(Kadavath et al., 2022). Consistency-based approaches cover SelfCheckGPT(Manakul et al., 2023), Lexical Similarity(Lin et al., 2022b), FActScore(Min et al., 2023) and RACE(Wang et al., 2025). Internal-state methods are represented by Inside(Chen et al., 2024a) and MIND(Su et al., 2024).
@@
Evaluation Metrics. We evaluate hallucination detection ability under two regimes following Janiak et al. (2025): ROUGE-based reference evaluation ($_{r}$) and LLM-as-a-Judge ($_{\mathrm{llm}}$). For performance measures, we report the area under the receiver operating characteristic curve (AUROC) and the area under the precision–recall curve (AUPRC). AUROC is widely used to assess the quality of binary classifiers and uncertainty estimators, while AUPRC highlights performance under class imbalance. In both cases, higher values indicate better detection.
@@ 4.2 Main Results
Q2: How does HalluGuard perform across LLMs of different scales? We further investigate whether the effectiveness of HalluGuard depends on model scale, as smaller backbones are typically more prone to hallucination. Table 3 reports representative results on small(Llama2-7B, Llama3-8B), mid-sized(Llama2-13B), and large-scale(Llama2-70B) models using SQuAD, GSM8K, and HaluEval. Across all settings, HalluGuard consistently surpasses baselines, with the largest margins on smaller models—for instance,
@@ 4.4 Case Study
Fine-grained hallucinations—lexically similar yet semantically incorrect outputs—pose a particular challenge for detection. To evaluate whether HalluGuard can comprehensively capture such subtle errors, we use the PAWS dataset(Zhang et al., 2019), which contrasts paraphrases with high surface overlap but divergent meanings. Following Li et al. (2025), we adopt ROUGE-based reference signals for evaluation (LABEL:tab:semantic_results). Across model scales, HalluGuard consistently surpasses baselines: it achieves 90.18% AUROC and 87.64% AUPRC on Llama2-70B, and 91.24% AUROC and 88.53% AUPRC on QwQ-32B—exceeding the next-best method by nearly five points. Even on GPT-2, it leads with 83.27% AUROC and 80.46% AUPRC. These results confirm HalluGuard’s effectiveness in capturing fine-grained semantic inconsistencies beyond benchmark settings.
@@ 5 Related Work
Detecting Data-Driven Hallucinations. Recent work has shown that internal activations encode rich indicators of such flaws. Chen et al. (2024a) proposed Eigenscore, which computes statistics of hidden representations from the eigen matrix to estimate hallucination risk. Su et al. (2024) introduced MIND, an unsupervised detector that models temporal dynamics of hidden states without requiring labels, along with HELM benchmark to enable standardized evaluation. Azaria and Mitchell (2023) demonstrated using linear probes on intermediate states to predict truthfulness.
Detecting Reasoning-Driven Hallucinations. There are other works targeting inference-time inconsistencies during generation—such as logical errors, instability across decoding steps, or temporal drift in extended outputs. Manakul et al. (2023) proposed SelfCheckGPT, which assesses self-consistency by sampling multiple candidate generations and measuring their alignment using entailment and lexical overlap. Kalai and Vempala (2024) introduced a suite of calibration-based uncertainty scores designed to capture hallucination risk directly from output distributions. Ding et al. (2025) proposed ReActScore, which integrates entropy with intermediate reasoning traces to detect failures in multi-step decision-making. FActScore(Min et al., 2023) decomposes outputs into atomic factual units and verifies each against retrieved passages using entailment-based scoring.
@@ 6 Conclusion
The reliability of LLMs is often undermined by hallucinations, which arise from two main sources: data-driven, caused by flawed knowledge acquired during training, and reasoning-driven, stemming from inference-time instabilities in multi-step generation. Although these hallucinations frequently evolve in practice, existing detectors usually target only one source and lack a solid theoretical foundation. To address this gap, we propose a unified theoretical framework–a Hallucination Risk Bound, which formally decomposes hallucination risk into data-driven and reasoning-driven components, offering a principled view of how hallucinations emerge and evolve during generation. Building on this foundation, we introduce HalluGuard, a NTK–based score that measures sensitivity to semantic perturbations and captures internal instabilities, thereby enabling holistic detection of both data-driven and reasoning-driven hallucinations. We evaluate HalluGuard across 10 diverse benchmarks, 11 competitive baselines, and 9 popular LLM backbones, where it consistently achieves state-of-the-art performance, demonstrating robustness and practical efficacy. Looking forward, leveraging HalluGuard’s sensitivity to error propagation offers a promising pathway for developing prognostic indicators in interactive multi-turn dialogues, enabling systems to predict and preempt hallucinations before they fully manifest.
@@ A.1 Assumptions Validation
Assumption A1 (Hilbert/RKHS structure with bounded second moment).
This assumption aligns with the classical Neural Tangent Kernel (NTK) approximation regime, where the model’s feature mapping is embedded in a reproducing kernel Hilbert space (RKHS) and the induced kernel admits a well-defined second moment. Such conditions are fundamental to the convergence and generalization analyses of infinitely wide neural networks, and are widely adopted in NTK theory (Jacot et al., 2020). In practice, bounded second-moment behavior is consistent with the hidden-state distributions observed across all evaluated LLMs, as reflected by stable activation statistics and NTK spectral profiles(Lee et al., 2020b).
Assumption A2 (Local Lipschitz continuity of the encoder $\Phi$).
This assumption reflects standard smoothness conditions in high-dimensional learning theory, ensuring that small perturbations in the input space induce controlled deviations in the encoded representation (Vershynin, 2018). Such local Lipschitz behavior is commonly invoked to guarantee stability under perturbations and is consistent with theoretical analyses of deep representations.
Assumption A3 (Local smoothness / second-order expansion).
This assumption corresponds to the classical NTK linearization framework, which approximates the behavior of wide neural networks through a local second-order expansion around a set of reference points (Lee et al., 2020a; Chizat et al., 2020). Importantly, our formulation requires this condition only locally around the $K$ sampled trajectories used by HalluGuard, rather than globally across the entire model parameter space. This localized validity preserves theoretical soundness while avoiding unrealistic global smoothness requirements that are known to be overly restrictive in large-scale models.
A.2 Bound Proof
We restate the main inequality from Section 3.2:
$\|u^{*}-u_{h}\|\leq\left[1+k_{\mathrm{pt}}\log\mathcal{O}(P,L)+k\frac{\epsilon_{\mathrm{mismatch}}}{\mathrm{Signal}_{k}}\right]\inf_{u\in U_{h}}\|u^{*}-u\|+\left|\mathcal{L}\right|\exp\left(-\frac{K\epsilon^{2}}{C}\right)\alpha\left(e^{\beta T}-1\right).$ (8)
Step 1: Triangle inequality split.
We define the hallucination decomposition by writing:
$\|u^{*}-u_{h}\|=\|u^{*}-\mathbb{E}[u_{h}]+\mathbb{E}[u_{h}]-u_{h}\|\leq\|u^{*}-\mathbb{E}[u_{h}]\|+\|u_{h}-\mathbb{E}[u_{h}]\|.$
We denote the first term as the deterministic approximation error (bias) and the second term as the stochastic residual (variance).
@@ A.2 Bound Proof
Assume $\mathbb{E}[u_{h}]$ is the Galerkin projection of $u^{*}$ in a coercive bilinear form $a(\cdot,\cdot)$, i.e., for all $v\in U_{h}$,
$a(\mathbb{E}[u_{h}],v)=\ell(v).$
Then, by Céa’s lemma, we have:
$\|u^{*}-\mathbb{E}[u_{h}]\|\leq\frac{\Lambda}{\gamma}\inf_{u\in U_{h}}\|u^{*}-u\|,$
where $\Lambda$ and $\gamma$ are continuity and coercivity constants of $a(\cdot,\cdot)$, respectively.
@@
Let $\ell_{h}:=\frac{1}{|\mathcal{L}|}\sum_{i=1}^{|\mathcal{L}|}\ell_{i}$ be the empirical supervision functional from finite labeled chains. Define the fluctuation:
$\Delta\ell:=\ell_{h}-\ell,$
and the residual:
$r:=u_{h}-\mathbb{E}[u_{h}],\quad\text{so that}\quad A_{h}r=\Delta\ell.$
Applying operator norm bounds and covering number uniformization (cf. Vershynin, 2018), we have with high probability:
$\|r\|\leq|\mathcal{L}|\exp\left(-\frac{K\epsilon^{2}}{C}\right)\alpha(e^{\beta T}-1),$
which completes the proof.
@@
Combining both terms yields:
@@
We now bound $\Lambda/\gamma$ via NTK decomposition.
A.3 Decomposition of NTK Continuity Constant
Let $a(\cdot,\cdot)$ denote the bilinear form induced by the NTK in the finite-width regime. We decompose:
@@ A.3 Decomposition of NTK Continuity Constant
where $a_{0}$ is the infinite-width baseline kernel, $\delta_{\mathrm{pt}}$ is the perturbation due to pre-training noise, and $\delta_{\mathrm{mm}}$ is the domain mismatch from fine-tuning. The continuity constant satisfies:
$\Lambda=\Lambda_{0}+\Delta_{\mathrm{pt}}+\Delta_{\mathrm{mm}}.$
Bounding $\Delta_{\mathrm{pt}}$.
Following Jacot et al. (2020), we apply matrix concentration to finite-width NTK:
@@
Bounding $\Delta_{\mathrm{mm}}$.
Using spectral generalization bounds under data distribution shift (Lee et al., 2020b), we have:
@@
Substituting both into the bound for $\Lambda/\gamma$, we get:

### v2

@@ 3.3 Hallucination Quantification via HalluGuard
Let $\mathcal{K}$ denote the NTK Gram matrix with eigenvalues $\lambda_{1}\geq\cdots\geq\lambda_{r}>0$ and condition number $\kappa(\mathcal{K})=\lambda_{\max}/\lambda_{\min}$. Let $J_{t}$ be the step-$t$ input-output Jacobian of the decoder, and define $\sigma_{\max}:=\sup_{t}\|J_{t}\|_{2}$ as the uniform spectral bound(note that $\sigma_{\max}$ is independent of the spectrum of $\mathcal{K}$).
Under Assumptions A1-A3, a standard NTK approximation argument yields $\inf_{u\in U_{h}}\|u^{*}-u\|\;\leq\;C_{d}\,\det(\mathcal{K})^{-c_{d}}\,\|u^{*}\|,$ so that $\det(\mathcal{K})$ capture the representations in systematic bias.
@@
In summary, $\det(\mathcal{K})$ quantifies representational adequacy, $\log\sigma_{\max}$ captures rollout amplification, and $\log\kappa^{2}$ penalizes spectral instability, together forming a compact and tractable proxy consistent with the Hallucination Risk Bound. The lightweight projection layers are self-supervised spectral calibration modules, optimized offline (via AdamW) to align NTK spectral properties across heterogeneous backbones into a stable, comparable geometric space-without hallucination labels or task-specific supervision, with the backbone fully frozen and zero runtime overhead during inference. Detailed proofs are provided in Appendix B.
@@ 4.1 Evaluation Setup
Benchmarks. We evaluate across 10 widely used benchmarks spanning three distinct categories. For data-grounded QA, we include RAGTruth (Niu et al., 2024), NQ-Open (Kwiatkowski et al., 2019), HotpotQA (Yang et al., 2018) and SQuAD (Rajpurkar et al., 2016), which emphasize factual correctness through external evidence. For reasoning-oriented tasks, we use GSM8K (Cobbe et al., 2021), MATH-500 (Hendrycks et al., 2021), and BBH (Suzgun et al., 2023), which require multi-step derivations prone to compounding errors. Finally, for instruction-following settings, we consider TruthfulQA (Lin et al., 2022a), HaluEval (Li et al., 2023a) and Natural (Wang et al., 2022), which probe hallucinations under open-ended or adversarial prompts.
Baselines. We compare HalluGuard with 11 competitive detectors spanning diverse strategies. Uncertainty-based methods include Perplexity (Ren et al., 2023), Length-Normalized Predictive Entropy(LN-Entropy) (Malinin and Gales, 2021), Semantic Entropy (Kuhn et al., 2023), Energy Score (Liu et al., 2020) and P(true) (Kadavath et al., 2022). Consistency-based approaches cover SelfCheckGPT (Manakul et al., 2023), Lexical Similarity (Lin et al., 2022b), FActScore (Min et al., 2023) and RACE (Wang et al., 2025). Internal-state methods are represented by Inside (Chen et al., 2024) and MIND (Su et al., 2024).
@@
Evaluation Metrics. We evaluate hallucination detection ability under two regimes following Janiak et al. (2025): ROUGE-based reference evaluation ($_{r}$) and LLM-as-a-Judge ($_{\mathrm{llm}}$). For performance measures, we report the area under the receiver operating characteristic curve (AUROC) and the area under the precision-recall curve (AUPRC). AUROC is widely used to assess the quality of binary classifiers and uncertainty estimators, while AUPRC highlights performance under class imbalance. In both cases, higher values indicate better detection.
@@ 4.2 Main Results
Q2: How does HalluGuard perform across LLMs of different scales? We further investigate whether the effectiveness of HalluGuard depends on model scale, as smaller backbones are typically more prone to hallucination. Table 3 reports representative results on small(Llama2-7B, Llama3-8B), mid-sized(Llama2-13B), and large-scale(Llama2-70B) models using SQuAD, GSM8K, and HaluEval. Across all settings, HalluGuard consistently surpasses baselines, with the largest margins on smaller models-for instance,
@@ 4.4 Case Study
Fine-grained hallucinations-lexically similar yet semantically incorrect outputs-pose a particular challenge for detection. To evaluate whether HalluGuard can comprehensively capture such subtle errors, we use the PAWS dataset (Zhang et al., 2019), which contrasts paraphrases with high surface overlap but divergent meanings. Following Li et al. (2025), we adopt ROUGE-based reference signals for evaluation (LABEL:tab:semantic_results). Across model scales, HalluGuard consistently surpasses baselines: it achieves 90.18% AUROC and 87.64% AUPRC on Llama2-70B, and 91.24% AUROC and 88.53% AUPRC on QwQ-32B-exceeding the next-best method by nearly five points. Even on GPT-2, it leads with 83.27% AUROC and 80.46% AUPRC. These results confirm HalluGuard’s effectiveness in capturing fine-grained semantic inconsistencies beyond benchmark settings.
@@ 5 Related Work
Detecting Data-Driven Hallucinations. Recent work has shown that internal activations encode rich indicators of such flaws. Chen et al. (2024) proposed Eigenscore, which computes statistics of hidden representations from the eigen matrix to estimate hallucination risk. Su et al. (2024) introduced MIND, an unsupervised detector that models temporal dynamics of hidden states without requiring labels, along with HELM benchmark to enable standardized evaluation. Azaria and Mitchell (2023) demonstrated using linear probes on intermediate states to predict truthfulness.
Detecting Reasoning-Driven Hallucinations. There are other works targeting inference-time inconsistencies during generation-such as logical errors, instability across decoding steps, or temporal drift in extended outputs. Manakul et al. (2023) proposed SelfCheckGPT, which assesses self-consistency by sampling multiple candidate generations and measuring their alignment using entailment and lexical overlap. Kalai and Vempala (2024) introduced a suite of calibration-based uncertainty scores designed to capture hallucination risk directly from output distributions. Ding et al. (2025) proposed ReActScore, which integrates entropy with intermediate reasoning traces to detect failures in multi-step decision-making. FActScore (Min et al., 2023) decomposes outputs into atomic factual units and verifies each against retrieved passages using entailment-based scoring.
@@ 6 Conclusion
The reliability of LLMs is often undermined by hallucinations, which arise from two main sources: data-driven, caused by flawed knowledge acquired during training, and reasoning-driven, stemming from inference-time instabilities in multi-step generation. Although these hallucinations frequently evolve in practice, existing detectors usually target only one source and lack a solid theoretical foundation. To address this gap, we propose a unified theoretical framework–a Hallucination Risk Bound, which formally decomposes hallucination risk into data-driven and reasoning-driven components, offering a principled view of how hallucinations emerge and evolve during generation. Building on this foundation, we introduce HalluGuard, a NTK-based score that measures sensitivity to semantic perturbations and captures internal instabilities, thereby enabling holistic detection of both data-driven and reasoning-driven hallucinations. We evaluate HalluGuard across 10 diverse benchmarks, 11 competitive baselines, and 9 popular LLM backbones, where it consistently achieves state-of-the-art performance, demonstrating robustness and practical efficacy. Looking forward, leveraging HalluGuard’s sensitivity to error propagation offers a promising pathway for developing prognostic indicators in interactive multi-turn dialogues, enabling systems to predict and preempt hallucinations before they fully manifest.
@@
Acknowledgements
We thank the anonymous reviewers for their constructive comments. This work is supported by the National Science Foundation under Award No. IIS-2339989 and No. 2406439, DARPA under contract No. HR00112490370 and No. HR001124S0013, U.S. Department of Homeland Security under Grant Award No. 17STCIN00001-08-00, Amazon-Virginia Tech Initiative for Efficient and Robust Machine Learning, Amazon AWS, Google, Cisco, 4-VA, Commonwealth Cyber Initiative, National Surface Transportation Safety Center for Excellence, and Virginia Tech. The views and conclusions are those of the authors and should not be interpreted as representing the official policies of the funding agencies or the government.
@@ A.1 Assumptions Validation
Setup For completeness, we briefly recall the main notation used in Section 3.2. Let $\mathcal{Y}$ denote the discrete metric space of finite-length token sequences. Let $U_{h}\subseteq\mathbb{R}^{d_{h}}$ be a $d_{h}$-dimensional Hilbert space equipped with inner product $\langle\cdot,\cdot\rangle$ and induced norm $\|\cdot\|$. The task-specific encoder $\Phi:\mathcal{Y}\to U_{h}$ is assumed to be $L_{\Phi}$-Lipschitz with respect to $d_{\mathcal{Y}}$.
Given input $\mathbf{x}$, the model defines a decoding distribution $P_{\theta}(\cdot\mid\mathbf{x})$ over $\mathcal{Y}$, and we denote the embedded random variable by $u_{h}:=\Phi(Y)$ where $Y\sim P_{\theta}(\cdot\mid\mathbf{x})$. For perturbations $\delta\in\mathbb{R}^{r}$ restricted to the local ball $\mathcal{B}_{\rho}$, the perturbed decoding distribution is denoted $P_{\theta}(\cdot\mid\mathbf{x},\delta)$, and the mean semantic response map is defined by $G_{Y}(\delta):=\mathbb{E}_{Y\sim P_{\theta}(\cdot\mid\mathbf{x},\delta)}[\Phi(Y)],$ with Jacobian $J=DG_{Y}(0)\in\mathbb{R}^{d_{h}\times r}$. The NTK Gram matrix on embedded perturbations is denoted $\mathcal{K}\in\mathbb{R}^{r\times r}$, with eigenvalues $\lambda_{1}\geq\dots\geq\lambda_{r}>0$ and condition number $\kappa(\mathcal{K})=\lambda_{\max}/\lambda_{\min}$. All expectations are taken with respect to the specified decoding distribution, and all norms are Euclidean unless otherwise stated.
Assumption A1 (Integrability and well-defined expectation). Assumption A1 ensures that the semantic embedding $\mathbb{E}_{Y\sim p_{\theta}(\cdot\mid\mathbf{x})}[\Phi(Y)]$ is well-defined as a Bochner expectation in the finite-dimensional Hilbert space $U_{h}$. The bounded second-moment condition guarantees that the expectation exists and is finite, which is a standard minimal requirement in stochastic analyses of neural network outputs. Such integrability assumptions are commonly adopted in NTK-based analyses (Jacot et al., 2018; Lee et al., 2020b), where control of second moments ensures stability of kernel spectra and well-posedness of linearized approximations.
Assumption A2 (Lipschitz continuity of the encoder $\Phi$). Assumption A2 imposes a controlled relationship between the discrete sequence space $(\mathcal{Y},d_{\mathcal{Y}})$ and the continuous embedding space $U_{h}$. The $L_{\Phi}$-Lipschitz condition ensures that bounded perturbations in edit distance induce proportionally bounded deviations in semantic representation. Such Lipschitz regularity is standard in high-dimensional learning theory (Vershynin, 2018) and is frequently invoked to establish stability under structured perturbations in representation learning. Importantly, this assumption is imposed only on the encoder map $\Phi$, not on the full autoregressive model.
Assumption A3 (Local Fréchet smoothness of the mean semantic response). Assumption A3 formalizes the local linearization principle underlying NTK theory. By requiring twice Fréchet differentiability of the mean response map $G_{Y}(\delta)=\mathbb{E}_{Y\sim P_{\theta}(\cdot\mid\mathbf{x},\delta)}[\Phi(Y)]$ within the perturbation ball $\mathcal{B}_{\rho}$, we ensure that $G_{Y}$ admits a controlled second-order expansion with uniform curvature constant $H_{\star}$. This local quadratic remainder bound is consistent with classical finite-width NTK linearization results (Lee et al., 2020a; Chizat et al., 2019), while avoiding unrealistic global smoothness requirements. Crucially, the assumption is imposed only on the expected semantic response, not on the discrete decoding distribution itself.
Remark.
Collectively, these assumptions provide a bridge between discrete autoregressive generation and continuous functional analysis. By restricting smoothness and curvature requirements to the localized perturbation neighborhood $\mathcal{B}_{\rho}$ and to the expectation-level map $G_{Y}$, we avoid imposing global regularity conditions over the infinite token space $\mathcal{Y}$. This localization ensures that the Hallucination Risk Bound is derived under mathematically controlled conditions while remaining aligned with the practical inference dynamics of large-scale language models.
A.2 Proof of Section 3.2
We restate the main inequality from Section 3.2. Note that due to the stochastic nature of autoregressive decoding, the bound holds with high probability. With probability at least $1-\delta$ over the generation process, the total hallucination risk satisfies:
[…]
Step 1: Triangle inequality split (Bias-Variance Decomposition).
Let $\bar{u}:=\mathbb{E}[u_{h}]$ be the expected semantic representation under the decoding distribution. By the triangle inequality in $U_{h}$, we decompose the hallucination risk into approximation error (bias) and stochastic residual (variance):
$\|u^{*}-u_{h}\|=\|u^{*}-\bar{u}+\bar{u}-u_{h}\|\leq\|u^{*}-\bar{u}\|+\|u_{h}-\bar{u}\|.$
@@
To bound the deterministic approximation error, we cast the model’s expected representation $\bar{u}$ as the solution to a variational problem in the Hilbert space $U_{h}$. Let $a(u,v):=\langle u,\mathcal{K}v\rangle_{U_{h}}$ denote the coercive bilinear form induced by the Neural Tangent Kernel (NTK) Gram operator $\mathcal{K}$, and let $f(v):=\langle u^{*},\mathcal{K}v\rangle_{U_{h}}$ be the bounded linear functional defining the target projection. Assuming $\bar{u}$ acts as the Galerkin projection of the target $u^{*}$ onto the trainable hypothesis space, it satisfies the weak formulation $a(\bar{u},v)=f(v)$ for all $v\in U_{h}$. By Céa’s lemma, the projection error is bounded by:
$\|u^{*}-\bar{u}\|\leq\frac{\Lambda}{\gamma}\inf_{u\in U_{h}}\|u^{*}-u\|,$
where $\Lambda$ and $\gamma$ are the continuity and coercivity constants of the NTK-induced bilinear form $a(\cdot,\cdot)$, respectively.
@@
We now bound the stochastic residual $\|u_{h}-\bar{u}\|$. Let $\mathcal{L}$ denote the set of $K$ independent sampled reasoning trajectories used during decoding. Under our local perturbation assumption (Assumption A3), the deviations of the hidden states are bounded by the local neighborhood radius $\rho$. Applying Bernstein’s inequality for bounded random vectors in a Hilbert space (Vershynin, 2018), the tail probabilities decay exponentially. For an error tolerance $\epsilon$ and an absolute constant $C>0$, we have with probability at least $1-\delta$:
$\|u_{h}-\bar{u}\|\leq|\mathcal{L}|\exp\left(-\frac{K\epsilon^{2}}{C}\right)\alpha(e^{\beta T}-1),$
where $\alpha$ is a scaling constant, $T$ is the sequence length, and $\beta\leq\log\sigma_{\max}$ bounds the per-step Jacobian spectral norm.
@@
Combining both terms yields the high-probability bound:
@@
We now bound the condition ratio $\Lambda/\gamma$ via NTK decomposition.
Step 5: Decomposition of NTK Continuity Constant
We decompose the bilinear form $a(\cdot,\cdot)$ into three components:
@@
where $a_{0}$ is the infinite-width baseline kernel, $\delta_{\mathrm{pt}}$ is the perturbation due to pre-training noise, and $\delta_{\mathrm{mm}}$ is the domain mismatch from fine-tuning. Consequently, the continuity constant satisfies $\Lambda=\Lambda_{0}+\Delta_{\mathrm{pt}}+\Delta_{\mathrm{mm}}$.
Bounding $\Delta_{\mathrm{pt}}$: Following standard matrix concentration bounds for finite-width NTKs (Jacot et al., 2018), the pre-training deviation scales logarithmically with the network parameters. Let $P$ be the number of parameters, $L$ the prompt length, and $k_{\mathrm{pt}}$ a pre-training scaling constant; we have:
@@
Bounding $\Delta_{\mathrm{mm}}$: Using spectral generalization bounds under data distribution shift (Lee et al., 2020b), the mismatch penalty is governed by the task-specific signal strength $\mathrm{Signal}_{k}$, the empirical mismatch error $\epsilon_{\mathrm{mismatch}}$, and a scaling constant $k$:
@@
Substituting both inequalities into the ratio for $\Lambda/\gamma$, and normalizing $\Lambda_{0}/\gamma\approx 1$, we obtain: